OpenStudy (anonymous):

Use either direct or inverse proportions to answer the following. Three men build a wall in ten days. How long would it take five men??

Parth (parthkohli):

They have already built it :/

OpenStudy (anonymous):

huh??

Parth (parthkohli):

Why would they build it again when the wall is already built?

OpenStudy (anonymous):

Dunno mate.

Parth (parthkohli):

@Coolsector it'd be inverse actually.

OpenStudy (anonymous):

yes looks like i did something wrong

OpenStudy (anonymous):

How do you know it would be inverse?@ParthKohli

Parth (parthkohli):

Because 5 people would take lesser time.

Parth (parthkohli):

I'll explain..wait

OpenStudy (anonymous):

3=k/10, k=30??

OpenStudy (anonymous):

5=30/x, x=6????

Parth (parthkohli):

http://www.youtube.com/watch?v=AQFZuih2odo

Parth (parthkohli):

Correct :D

OpenStudy (anonymous):

so k=30 days 1 person

Parth (parthkohli):

Yes. Right!

OpenStudy (anonymous):

Thanks man.

Parth (parthkohli):

Heehee

OpenStudy (anonymous):

@ParthKohli how would I check my answer??

Parth (parthkohli):

Where is it?

OpenStudy (anonymous):

my answer of 6 days